# Bench: Extending Long Context Evaluation Beyond 100K Tokens

### 22 Feb 2024 v2

@@ Abstract
Processing and reasoning over long contexts is crucial for many practical applications of Large Language Models (LLMs), such as document comprehension and agent construction. Despite recent strides in making LLMs process contexts with more than 100K tokens, there is currently a lack of a standardized benchmark to evaluate this long-context capability. Existing public benchmarks typically focus on contexts around 10K tokens, limiting the assessment and comparison of LLMs in processing longer contexts. In this paper, we propose Bench, the first LLM benchmark featuring an average data length surpassing 100K tokens. Bench comprises synthetic and realistic tasks spanning diverse domains, presented in both English and Chinese. The tasks in Bench are designed to require well understanding of long dependencies in contexts, and make simply retrieving a limited number of passages from contexts not sufficient for these tasks. In our experiments, based on Bench, we evaluate the state-of-the-art proprietary and open-source LLMs tailored for processing long contexts. The results indicate that existing long context LLMs still require significant advancements to effectively process 100K+ context. We further present three intriguing analyses regarding the behavior of LLMs processing long context.
@@ 1 Introduction
Recent advancements in training infrastructure Shoeybi et al. (2019); Narayanan et al. (2021); Dao et al. (2022); Dao (2023), and efforts to improve length generalization Anil et al. (2022); Chen et al. (2023b); Peng et al. (2023b)111https://www.reddit.com/r/LocalLLaMA/comments/14lz7j5/ntkaware_scaled_rope_allows_llama_models_to_have/ have led to rapid developments in long-context LLMs. Based on these improved training infrastructures and length generalization methods, several LLMs have purportedly managed to process data exceeding 100K tokens (Peng et al., 2023b; OpenAI, 2023b; 01.AI, 2023b, a), with Claude 2 Anthropic (2023) and Kimi-Chat AI (2023) even claiming to be able to process up to 200K tokens. However, the rapid emergence of long-context LLMs has outpaced the development of adequate evaluation benchmarks. Present long-context benchmarks predominantly feature contexts averaging around 10K tokens (Bai et al., 2023; Tay et al., 2020), invariably falling below 100K tokens. This lag in the advancement of long-context evaluation methodologies impedes both the comparative analysis of diverse long-context LLMs and the pinpointing of potential enhancements in long-context processing.
@@ 2 Related Work
Recently, many LLMs have shown the ability to handle over 100K tokens. Some popular proprietary 100K+ LLMs include GPT-4, Claude 2 (Anthropic, 2023), and Kimi-Chat (AI, 2023). On the other hand, there are much fewer open-source 100K+ models. Some notable models include YaRN (Peng et al., 2023b) and Yi-200K (01.AI, 2023a, b). In this paper, we benchmark GPT-4, Claude 2, Kimi-Chat, and YaRN-Mistral-7B-128K222https://huggingface.co/NousResearch/Yarn-Mistral-7b-128k, we denote this model by YaRN-Mistral. on Bench, which are some of the latest and strongest LLMs that claim to be able to handle over 100K tokens.
@@ 3.1.1 Novel
We develop novel-based tasks as outlined in Figure 3, utilizing novels sourced from websites333https://www.sparknotes.com/444https://www.cliffsnotes.com/ and are manually filtered. More annotation information in Appendix. C.
@@ 3.1.2 Dialogue
The construction process for the En.Dia task is depicted in Figure 3. We gather movie and drama scripts from a designated online database555https://imsdb.com/, focusing on a corpus of long, multi-role dialogues. Only the English scripts are retained and necessary cleaning is applied.
@@ 3.1.3 Code
We develop the task as per the process illustrated in Figure 3. Code repositories, sourced from PyPI666https://pypi.org/, undergo a filtering process, and those outside the 64K to 256K token range are excluded (tokenization via the tiktoken tokenizerOpenAI (2023c)). Each repository is transformed into a single file, aggregating the content from all files within, each prefaced by its relative path to the root directory. Three of the authors then insert a deliberate and obvious error into one function per repository. The options are presented in the Class.Function format. Six methods are employed for bug insertion: (1) deleting a necessary variable declaration; (2) using an incorrect number of arguments in function calls; (3) creating infinite loops; (4) causing indentation errors; (5) substituting references with undefined variable/function names; (6) introducing blatant syntax errors (e.g., non-closed brackets).
@@ Appendix A RWKV
Table 4 shows the performance of RWKV-4-World-7 in comparison to our baselines. We find that RWKV-4-World-7B outputs unintelligible texts on our benchmark, which causes it to achieve zero performance on Retrieve.PassKey, which is the easiest task for other baselines. This is likely because this model was not trained on inputs of this length and suffers from train-test domain shift.777We emphasize that this result is not evidence that the architecture of RWKV is incapable of handling lengthy inputs. Therefore, we do not consider testing it on other tasks in our benchmark.
@@ Appendix B Prompt Templates
The prompt below is used by GPT-4888It should be noted that, when using other templates, GPT-4 has a strong tendency to reject to perform this task by claiming that such the platform is not designed to complete such tasks.:

### 24 Feb 2024 v3

@@ Abstract
Processing and reasoning over long contexts is crucial for many practical applications of Large Language Models (LLMs), such as document comprehension and agent construction. Despite recent strides in making LLMs process contexts with more than 100K tokens, there is currently a lack of a standardized benchmark to evaluate this long-context capability. Existing public benchmarks typically focus on contexts around 10K tokens, limiting the assessment and comparison of LLMs in processing longer contexts. In this paper, we propose Bench, the first LLM benchmark featuring an average data length surpassing 100K tokens. Bench comprises synthetic and realistic tasks spanning diverse domains, presented in both English and Chinese. The tasks in Bench are designed to require well understanding of long dependencies in contexts, and make simply retrieving a limited number of passages from contexts not sufficient for these tasks. In our experiments, based on Bench, we evaluate the state-of-the-art proprietary and open-source LLMs tailored for processing long contexts. The results indicate that existing long context LLMs still require significant advancements to effectively process 100K+ context. We further present three intriguing analyses regarding the behavior of LLMs processing long context. Our code and data is released111https://github.com/OpenBMB/InfiniteBench222https://huggingface.co/datasets/xinrongzhang2022/InfiniteBench.
@@ 1 Introduction
Recent advancements in training infrastructure Shoeybi et al. (2019); Narayanan et al. (2021); Dao et al. (2022); Dao (2023), and efforts to improve length generalization Anil et al. (2022); Chen et al. (2023b); Peng et al. (2023b)333https://www.reddit.com/r/LocalLLaMA/comments/14lz7j5/ntkaware_scaled_rope_allows_llama_models_to_have/ have led to rapid developments in long-context LLMs. Based on these improved training infrastructures and length generalization methods, several LLMs have purportedly managed to process data exceeding 100K tokens (Peng et al., 2023b; OpenAI, 2023b; 01.AI, 2023b, a), with Claude 2 Anthropic (2023) and Kimi-Chat AI (2023) even claiming to be able to process up to 200K tokens. However, the rapid emergence of long-context LLMs has outpaced the development of adequate evaluation benchmarks. Present long-context benchmarks predominantly feature contexts averaging around 10K tokens (Bai et al., 2023; Tay et al., 2020), invariably falling below 100K tokens. This lag in the advancement of long-context evaluation methodologies impedes both the comparative analysis of diverse long-context LLMs and the pinpointing of potential enhancements in long-context processing.
@@ 2 Related Work
Recently, many LLMs have shown the ability to handle over 100K tokens. Some popular proprietary 100K+ LLMs include GPT-4, Claude 2 (Anthropic, 2023), and Kimi-Chat (AI, 2023). On the other hand, there are much fewer open-source 100K+ models. Some notable models include YaRN (Peng et al., 2023b) and Yi-200K (01.AI, 2023a, b). In this paper, we benchmark GPT-4, Claude 2, Kimi-Chat, and YaRN-Mistral-7B-128K444https://huggingface.co/NousResearch/Yarn-Mistral-7b-128k, we denote this model by YaRN-Mistral. on Bench, which are some of the latest and strongest LLMs that claim to be able to handle over 100K tokens.
@@ 3.1.1 Novel
We develop novel-based tasks as outlined in Figure 3, utilizing novels sourced from websites555https://www.sparknotes.com/666https://www.cliffsnotes.com/ and are manually filtered. More annotation information in Appendix. C.
@@ 3.1.2 Dialogue
The construction process for the En.Dia task is depicted in Figure 3. We gather movie and drama scripts from a designated online database777https://imsdb.com/, focusing on a corpus of long, multi-role dialogues. Only the English scripts are retained and necessary cleaning is applied.
@@ 3.1.3 Code
We develop the task as per the process illustrated in Figure 3. Code repositories, sourced from PyPI888https://pypi.org/, undergo a filtering process, and those outside the 64K to 256K token range are excluded (tokenization via the tiktoken tokenizerOpenAI (2023c)). Each repository is transformed into a single file, aggregating the content from all files within, each prefaced by its relative path to the root directory. Three of the authors then insert a deliberate and obvious error into one function per repository. The options are presented in the Class.Function format. Six methods are employed for bug insertion: (1) deleting a necessary variable declaration; (2) using an incorrect number of arguments in function calls; (3) creating infinite loops; (4) causing indentation errors; (5) substituting references with undefined variable/function names; (6) introducing blatant syntax errors (e.g., non-closed brackets).
@@ Appendix A RWKV
Table 4 shows the performance of RWKV-4-World-7 in comparison to our baselines. We find that RWKV-4-World-7B outputs unintelligible texts on our benchmark, which causes it to achieve zero performance on Retrieve.PassKey, which is the easiest task for other baselines. This is likely because this model was not trained on inputs of this length and suffers from train-test domain shift.999We emphasize that this result is not evidence that the architecture of RWKV is incapable of handling lengthy inputs. Therefore, we do not consider testing it on other tasks in our benchmark.
@@ Appendix B Prompt Templates
The prompt below is used by GPT-4101010It should be noted that, when using other templates, GPT-4 has a strong tendency to reject to perform this task by claiming that such the platform is not designed to complete such tasks.: